gababb09:

If (5^0)^x=1, what are the possible values of x? explain your answer.

Laylalyssa:

ok so on this we have the following expression: (5^0)^x1=1

Laylalyssa:

for the def of power properties we have to: (a^n)^m=a^n*m

Laylalyssa:

After this we re-write

Laylalyssa:

can u rewrite it

Laylalyssa:

?

gababb09:

hang on

Laylalyssa:

ight

SemiDefinite:

what is \(5^0\) ?

gababb09:

5^0=1

gababb09:

Any number with an exponent of zero is equal to 1.